If (1 + x) (1 + x + x2) (1 + x + x2 + x3) ...... (1 + x + x2 + x3 + ...... + xn) a0 + a1x + a2x2 + a3x3 + ...... + amxm then has the value equal to
n!
(n + 1) !
(n - 1)!
none

Grade:12th Pass

1 Answers

Nehal Wani
13 years ago

  =a0 + a1(1) + a2(1)2 + a3(1)3 + ...... + am(1)m

 =a0 + a1x + a2x2 + a3x3 + ...... + amxm ………where x=1

=(1 + x) (1 + x + x2) (1 + x + x2 + x3) ...... (1 + x + x2 + x3 + ...... + xn)

 =1.2.3.4.5.6……………………………n=n!